an isosceles triangel has a perimeter of 34cm. the congurant mesaure (4x-1) cm. the thrid leg is x. what is the lenght of each side?
Answer by mananth(16949)

perimeter = sum of all sides
(4x-1)(4x-1)+x=34
8x-2+x=34
9x-2=34
add 2
9x=34+2
9x=36
/9
x=4
...
congruent sides = 4x-1
16-1 = 15
...
15+15+4 = 34
...
m.ananth@hotmail.ca